# Blinded by Generated Contexts: How Language Models Merge Generated and Retrieved Contexts for Open-Domain QA?

### v4

@@ Abstract
While auxiliary information has become a key to enhancing Large Language Models (LLMs), relatively little is known about how LLMs merge these contexts, specifically contexts generated by LLMs and those retrieved from external sources. To investigate this, we formulate a systematic framework to identify whether LLMs’ responses, derived from the integration of generated and retrieved contexts, are attributed to either generated or retrieved contexts. To easily trace the origin of the response, we construct datasets with conflicting contexts, i.e., each question is paired with both generated and retrieved contexts, yet only one of them contains the correct answer. Our experiments reveal a significant bias in several LLMs (GPT-4/3.5 and Llama2) to favor generated contexts, even when they provide incorrect information. We further identify two key factors contributing to this bias: i) contexts generated by LLMs typically show greater similarity to the questions, increasing their likelihood of being selected; ii) the segmentation process used in retrieved contexts disrupts their completeness, thereby hindering their full utilization in LLMs. Our analysis enhances the understanding of how LLMs merge diverse contexts, offering valuable insights for advancing current augmentation methods for LLMs [^1] Code released at https://anonymous.4open.science/r/7BB7/.
@@
Merge Generated and Retrieved Contexts for Open-Domain QA?
Hexiang Tan♠♡, Fei Sun [^2] Corresponding author.  , Wanli Yang♠♢, Yuanzhuo Wang, Qi Cao, Xueqi Cheng♠♡ CAS Key Laboratory of AI Safety & Security, Institute of Computing Technology, Chinese Academy of Sciences, Beijing, China University of Chinese Academy of Sciences, Beijing, China Nankai University, Tianjin, China {tanhexiang21s, sunfei, wangyuanzhuo, caoqi, cxq}@ict.ac.cn
@@ 1 Introduction
Recent advancements in augmenting Large Language Models (LLMs) with auxiliary information have significantly revolutionized their efficacy in knowledge-intensive tasks Chang et al. (2023); Ram et al. (2023). In this evolving field, existing works can be broadly categorized into two groups based on information sources: generation-augmented and retrieval-augmented approaches. To effectively harness the internal knowledge of LLMs, generation-augmented approaches Liu et al. (2022); Sun et al. (2023), e.g., GenRead Yu et al. (2022), instruct LLMs to initially generate a background context tailored to the given question, which is then employed as the basis for producing the final answer. In contrast, retrieval-augmented approaches Lewis et al. (2020); Ram et al. (2023) adopt an alternative strategy by incorporating relevant passages from external corpora, e.g., Wikipedia, as context, thereby notably enhancing LLMs’ capability to address challenges like knowledge updates Jang et al. (2022) and long-tail knowledge Kandpal et al. (2023).
@@
Building on the foundations laid by generation-augmented and retrieval-augmented methods, recent hybrid approaches have attempted to integrate them to further improve performance in tasks like Question Answering (QA) Yu et al. (2022); Mallen et al. (2023). These hybrid approaches face a significant challenge: conflicts between diverse sources can impede the effectiveness of information integration Zhang et al. (2023). While recent works have investigated conflicts within contexts from a single source, either only retrieved (Chen et al., 2022) or generated (Xie et al., 2023), it remains unclear how LLMs resolve conflicts between generated and retrieved contexts. This study, therefore, aims to investigate the underlying mechanisms by which LLMs process the two types of contexts, especially when they contain conflicting information.
@@
Our investigation was driven by a striking observation: in certain cases, models relying solely on retrieval contexts succeeded, whereas, counter-intuitively, hybrid approaches failed, as depicted in Figure 1. To uncover the underlying reasons, we proposed a systematic framework to dissect the process by which LLMs merge generated and retrieved contexts. We curated tailored datasets in which each question is accompanied by a pair of generated and retrieved contexts. These contexts are deliberately designed to be inconsistent, with only one containing the correct answer to its corresponding question. These datasets provide a solid foundation for determining whether LLMs utilize retrieved or generated context to produce responses in QA tasks.
@@
In this paper, we conducted a series of controlled experiments using our uniquely designed datasets to empirically study this question, focusing on several state-of-the-art closed (GPT-3.5/4) and open (Llama2-7b/13b) LLMs. Surprisingly, our findings reveal a pronounced bias in LLMs to favor generated contexts, even when the generated contexts offer incorrect information while the retrieval contexts hold the correct answers. Furthermore, this bias persists regardless of whether the generated text was produced internally or by other LLMs (§4.2). These findings highlight a critical challenge for existing LLMs in effectively merging internal parametric knowledge (i.e., generated contexts) and external information (i.e., retrieved contexts), under increasingly common non-tunable settings, e.g., those involving black-box APIs like GPT-4.
@@
Furthermore, extensive controlled experiments are conducted to investigate the underlying causes of the bias and provide the following insights: (i) confirmation bias (Xie et al., 2023) is not a key factor5.1): LLMs maintain a significant preference for generated contexts when they contain information inconsistent with LLMs’ parametric knowledge. (ii) text similarity is a significant factor5.2): compared to retrieved contexts, generated contexts typically exhibit a higher degree of similarity to the questions, even when they contain incorrect information. The samples with a larger similarity gap between generated and retrieved contexts exhibit a more pronounced bias. These findings emphasize the need for caution with LLM-generated contexts, to avoid being misled by highly relevant but inaccurate information. (iii) semantic completeness matters5.3): LLMs tend to favor contexts with semantic integrity. The segmentation process used in retrieved contexts may disrupt their completeness, thereby hindering their full utilization in LLMs.
This work preliminarily explores the growing challenge of LLMs utilizing contexts from diverse sources, especially in light of the increasing prevalence of LLM-generated content on the internet, which may contain potential misinformation (Pan et al., 2023). Furthermore, our findings offer valuable guidance for enhancing existing retrieval-augmented methods, such as optimizing passage segmentation in retrieval systems. Our main contributions can be summarized as:
@@
  • To facilitate controlled experiments, we develop a specialized framework for constructing tailored datasets and excluding confounding factors, e.g., input order and context length.
  • Our extensive analyses have identified two key factors, i.e., text similarity and semantic completeness, in the context utilization of LLMs. Moreover, we reveal that the confirmation bias (Xie et al., 2023) cannot account for the bias in this paper.
@@ 2.1 Background
Knowledge Conflicts within Contexts. These studies mainly focus on conflicts within a single type of input contexts, either only retrieved (Chen et al., 2022) or generated (Xie et al., 2023), leaving underexplored how LLMs resolve conflicts between diverse contexts. A systematic review of related works is provided in Section A.1.
@@ 2.2 Answer Tracing Task
We specifically focus on non-tunable, i.e. zero-shot setting, LLMs acting as the reader and generator, reflecting prevalent real-world use cases like ChatGPT. This direction is motivated by the high cost and limited accessibility of fine-tuning, which makes the direct use of non-tunable LLMs popular. Additionally, given the extensive use of LLMs, any bias or issue in their merging mechanisms could lead to serious consequences.
@@ 3.1 Context-Conflicting Datasets
Context Preparation. Figure 4 Step 1 illustrates the process of preparing contexts for each question. For retrieved contexts, it is obtained from the top-$1$ ranked passage from Wikipedia using Contriever (Izacard et al., 2021), a powerful off-the-shelf retrieval model that is extensively employed in various retrieval-augmented generation systems Shi et al. (2023); Ram et al. (2023).
@@
Sample Filtering for Traceability. With each question paired with a single context (either generated or retrieved) established in the initial stage, the reader generates the corresponding candidate answer, as shown in Step 2 of Figure 4. To unravel the mechanisms of LLMs in context merging, it is essential to ensure the traceability, i.e., the output answer is derived from the input context, rather than the intrinsic parametric knowledge of the LLMs. To achieve this, we only keep samples in which both the generated and retrieved contexts exactly include their respective generated answers, exemplified by $a_{\phi}^{\gamma}{\in}d_{1}^{\gamma}$, where $\in$ denotes $d_{1}^{\gamma}$ contains the substring $a_{\phi}^{\gamma}$. This practice is grounded in the findings of Chen et al. (2022); Xie et al. (2023), which demonstrate that in the presence of external context, LLMs tend to rely on external context rather than their intrinsic parametric knowledge.
Building Context-Conflicting Dataset. Having obtained answers for each type of context, we are now positioned to construct our context-conflicting (CC) datasets, as depicted in Step 3 of Figure 4. Initially, We employ the exact match metric Yu et al. (2022) to evaluate the correctness of candidate answers derived from contexts, considering an answer correct if its normalized form matches any of the golden answers. Subsequently, the context-conflicting datasets are composed of samples for which only one of the two types of contexts, either generated or retrieved, yields the correct answer, thereby ensuring the exclusivity. Notably, each dataset comprises two distinct subsets: AIG, consisting of samples with correct answers only in the generated context; and AIR comprising samples with correct answers only in the retrieved context.
@@ 3.2 Statistics of Datasets
Table 1: Dataset statistics across LLMs, “Avg. Prop.” shows average proportions of subsets to original datasets.
@@
We initially adopt a typical and simple setting in which a singular LLM serves as both the generator and reader. Table 1 provides statistics for the constructed subsets corresponding to various LLMs, including GPT-4 (gpt-4-0613), GPT-3.5 (gpt-3.5-turbo-0613), Llama2-7b/13b (Llama-2-7b/13b-chat (Touvron et al., 2023)). The statistics show that the context-conflicting subsets form a substantial part of the datasets, underscoring the need to investigate how LLMs integrate these distinct contexts. Notably, GPT-4 has fewer conflicting instances than other LLMs, because of its higher efficacy in answering questions using either solely retrieved or generated contexts.
Section 4.2 also explores a more complex scenario in which the generator and reader are distinct LLMs and show the statistics in Section A.3.
@@ 3.3 Evaluation Metric
The metric $\mathrm{DiffGR}$, ranging from $[-1,1]$, quantifies the extent of LLMs’ tendency to rely on generated contexts over retrieved contexts. Using AIR as an example, where all correct answers come from retrieved contexts, an ideal $\mathrm{DiffGR}$ value should be $-1$, i.e., LLMs should always rely on retrieved contexts.
@@ 4 How LLMs Merge Contexts?
This section conducts experiments on the developed datasets to investigate the merging mechanism of the LLMs in hybrid approaches in two settings. We first consider a typical setting where the generator and reader share a single LLM, to explore how LLMs merge retrieved and self-generated contexts (§4.1). Then, we extend our experiments to more flexible combinations of generator and reader (§4.2) to investigate their effects.
@@ 4.1 LLMs Prefer Self-Generated Contexts
Our preliminary experiments, in which a single LLM serves both as generator and reader, are designed to explore how LLMs integrate information from retrieved and self-generated contexts. The LLMs under evaluation are tasked with answering questions using both types of contexts on their corresponding context-conflicting subsets. In all experiments, we employ a randomized input sequence of contexts to mitigate the influence of order, which is further discussed in Appendix A.4.
We begin our analysis by examining LLMs’ QA performance on context-conflicting datasets to reveal how well can LLMs utilize both types of contexts. Table 2 presents the Exact Match scores (Yu et al., 2022) across various LLMs. Surprisingly, LLMs demonstrate significantly low performance ($\leq 22.16\%$) on AIR subsets, despite the fact that the retrieved context alone consistently yields the correct answer on these subsets. In contrast, LLMs exhibit strong performance on AIG subsets (most near $90\%$). Overall, all LLMs exhibit a significant performance gap between AIR and AIG datasets, with a pronounced decline in performance when the correct answers come from retrieved contexts.
@@
To further reveal LLMs’ behavior underlying the QA performance, we trace the source contexts of LLMs’ answers using the proposed $\mathrm{DiffGR}$ metric. An ideal LLM should always rely on retrieved contexts on AIR subsets ($\mathrm{DiffGR=-1}$), and always rely on generated contexts on AIG subsets ($\mathrm{DiffGR=1}$). Contrary to expectations, Figure 5 illustrates that LLMs fail to identify the correct information and consistently tend to rely on generated contexts on both AIG and AIR subsets. This result indicates a pronounced bias in LLMs to favor generated contexts, even when they provide incorrect information. This bias leads to the insufficient utilization of retrieved contexts mentioned above and highlights a critical challenge for existing LLMs in effectively merging generated and retrieved contexts. As the bias on AIR subsets has a more direct impact on the performance, the following experiments and analysis will focus on the biases on these subsets to conserve space. Results on the AIG subsets can be found in Appendix A.5.
@@ 4.2 LLMs Broadly Prefer Generated Contexts
The above experiments reveal the bias in LLMs to favor the self-generated context. Consequently, an intriguing question emerges: Do they also exhibit a similar bias towards contexts generated by other LLMs? To investigate this question, this section extends the experiments to more flexible combinations of generators and readers. This setting is also of practical significance, as recent works have explored the decoupling of generators and readers to achieve modularization of knowledge (Luo et al., 2023; Feng et al., 2023).
We construct context-conflicting datasets for each (generator, reader) pair respectively, as detailed in Appendix A.3. Based on these datasets, we then compute $\mathrm{DiffGR}$ metric to examine biases across various (generator, reader) pairs, as shown in Figure 6, and observe two notable insights: (i) LLMs also biased towards contexts generated by other LLMs.This suggests that such bias in LLMs is widespread and not limited to self-generated contexts. (ii) LLMs usually exhibit a stronger bias to contexts generated by themselves.The sole exception is Llama2-7b, which shows the strongest bias when paired with Llama2-13b as its generator. This phenomenon likely results from their highly similar model structures and training processes (Touvron et al., 2023).

### v6

# Blinded by Generated Contexts: How Language Models Merge Generated and Retrieved Contexts When Knowledge Conflicts?

@@ Abstract
While auxiliary information has become a key to enhancing Large Language Models (LLMs), relatively little is known about how LLMs merge these contexts, specifically contexts generated by LLMs and those retrieved from external sources. To investigate this, we formulate a systematic framework to identify whether LLMs’ responses are attributed to either generated or retrieved contexts. To easily trace the origin of the response, we construct datasets with conflicting contexts, i.e., each question is paired with both generated and retrieved contexts, yet only one of them contains the correct answer. Our experiments reveal a significant bias in several LLMs (GPT-4/3.5 and Llama2) to favor generated contexts, even when they provide incorrect information. We further identify two key factors contributing to this bias: i) contexts generated by LLMs typically show greater similarity to the questions, increasing their likelihood of being selected; ii) the segmentation process used in retrieved contexts disrupts their completeness, thereby hindering their full utilization in LLMs. Our analysis enhances the understanding of how LLMs merge diverse contexts, offers valuable insights for advancing current LLM augmentation methods, and highlights the risk of generated misinformation for retrieval-augmented LLMs [^1] Code released at https://github.com/Tan-Hexiang/RetrieveOrGenerated..
@@
Merge Generated and Retrieved Contexts When Knowledge Conflicts?
Hexiang Tan♠♡, Fei Sun [^2] Corresponding author., Wanli Yang, Yuanzhuo Wang, Qi Cao, Xueqi Cheng♠♡ CAS Key Laboratory of AI Safety, Institute of Computing Technology, Chinese Academy of Sciences, Beijing, China University of Chinese Academy of Sciences, Beijing, China {tanhexiang21s, sunfei, wangyuanzhuo, caoqi, cxq}@ict.ac.cn
@@ 1 Introduction
Recent advancements in augmenting Large Language Models (LLMs) with auxiliary information have significantly revolutionized their efficacy in knowledge-intensive tasks Chang et al. (2023); Ram et al. (2023). This auxiliary information can originate from contexts generated by LLMs or retrieved from external sources. For the former, Liu et al. (2022); Sun et al. (2023); Yu et al. (2022) instruct LLMs to initially generate a background context tailored to the given question, which then serves as the basis for producing the final answer. In contrast, retrieval-augmented approaches (Lewis et al., 2020; Ram et al., 2023) include relevant passages retrieved from external corpora, such as Wikipedia, as context, thereby notably enhancing LLMs’ capability to address challenges like knowledge updates Jang et al. (2022) and long-tail knowledge Kandpal et al. (2023).
@@
Recent hybrid methods have attempted to integrate these two types of contexts to further improve performance in tasks like Question Answering (QA) (Yu et al., 2022; Mallen et al., 2023; Zhang et al., 2023). However, we observe an abnormal phenomenon: in certain cases, models relying solely on retrieval contexts succeeded, whereas hybrid approaches failed, as depicted in Figure 1. This observation implies that LLMs may struggle to effectively integrate diverse types of contexts, overlooking correct information in retrieved contexts. To uncover the reasoning behind this, this study investigates the underlying mechanisms by which LLMs process the two types of contexts, especially when they contain conflicting information.
@@
To facilitate the investigation, we proposed a systematic framework to dissect the process by which LLMs merge generated and retrieved contexts. We curated tailored context-conflicting (CC) datasets in which each question is accompanied by a pair of generated and retrieved contexts. These contexts are deliberately designed to be inconsistent, with only one containing the correct answer to its corresponding question. These datasets provide a solid foundation for determining whether LLMs utilize retrieved or generated context to produce responses in QA tasks.
@@
In this paper, we conducted a series of controlled experiments using our uniquely designed datasets to empirically study this question, focusing on several state-of-the-art closed (GPT-3.5/4) and open (Llama2-7b/13b) LLMs. Surprisingly, our findings reveal a pronounced bias in LLMs to favor generated contexts across various LLMs (§4.2), even when the generated contexts offer incorrect information while the retrieval contexts hold the correct answers. Further analysis shows that the bias is prevalent across various retrieval models (§4.3). These findings highlight a critical challenge for existing LLMs in effectively merging contexts from diverse sources, especially in light of the increasing prevalence of LLM-generated content on the internet, which may contain potential misinformation (Pan et al., 2023; Chen and Shu, 2023).
@@
Through extensive empirical analyses, we reveal several factors contributing to this bias: (i) confirmation bias (Xie et al., 2023) is not a key factor5.1): LLMs maintain a significant preference for generated contexts even when they are inconsistent with LLMs’ parametric knowledge. (ii) text similarity is a significant factor5.2): compared to retrieved contexts, generated contexts typically exhibit a higher degree of similarity to the questions, even when they contain incorrect information. The samples with a larger similarity gap between generated and retrieved contexts exhibit a more pronounced bias. (iii) semantic completeness matters5.3): LLMs tend to favor contexts with semantic integrity. The segmentation process used in retrieved contexts may disrupt their completeness, thereby hindering their full utilization in LLMs. This finding emphasizes the need for optimizing passage segmentation in current retrieval systems.
In summary, we explore the challenges LLMs face when utilizing conflicting contexts and make the following contributions:
@@
  • To facilitate controlled experiments, we develop a specialized framework for constructing tailored datasets and excluding confounding factors, e.g. input order, and context length.
  • Our extensive analyses have identified two key factors, i.e., text similarity and semantic completeness, in the context utilization of LLMs.
@@ 2.1 Background
Knowledge Conflicts within Contexts. These studies mainly focus on conflicts within a single type of input contexts, either only retrieved (Chen et al., 2022) or generated (Xie et al., 2023), leaving underexplored how LLMs resolve conflicts between diverse contexts.
@@ 2.2 Answer Tracing Task
We specifically focus on non-tunable LLMs, i.e. in zero-shot settings, to reflect prevalent real-world use cases like ChatGPT. This direction is motivated by the high cost and limited accessibility of fine-tuning, which makes the direct use of non-tunable LLMs popular. Additionally, given the extensive use of LLMs, any bias or issue in their merging mechanisms could lead to serious consequences.
@@ 3.1 Context-Conflicting Datasets
Context Preparation. Step 1 in Figure 4 illustrates the process of preparing contexts for each question. For retrieved contexts, it is obtained from the top-$1$ ranked passage from Wikipedia using Contriever (Izacard et al., 2021), a powerful off-the-shelf retrieval model that is extensively employed in various retrieval-augmented generation systems Shi et al. (2023); Ram et al. (2023).
@@
Sample Filtering for Traceability. With each question paired with a single context (either generated or retrieved) established in the initial stage, the reader generates the corresponding candidate answer, as shown in Step 2 of Figure 4. To unravel the mechanisms of LLMs in context merging, it is essential to ensure the traceability, i.e., the output answer is derived from the input context, rather than the intrinsic parametric knowledge of the LLMs. To achieve this, we only keep samples in which both the generated and retrieved contexts exactly include their respective generated answers, exemplified by $a_{\phi}^{\gamma}{\in}d_{1}^{\gamma}$, where $\in$ denotes $d_{1}^{\gamma}$ contains the answer string $a_{\phi}^{\gamma}$. This practice is grounded in the findings of Chen et al. (2022); Xie et al. (2023), which demonstrate that in the presence of external context, LLMs mostly rely on external context rather than their intrinsic parametric knowledge. Despite these efforts, the complete elimination of the influence from parametric knowledge remains challenging. Therefore, we separately examine its impact in §5.1, where we demonstrate that it has a negligible effect on our conclusions.
Building Context-Conflicting Dataset. Having obtained answers for each type of context, we are now positioned to construct our context-conflicting (CC) datasets, as depicted in Step 3 of Figure 4. Initially, We employ the exact match metric Yu et al. (2022) to evaluate the correctness of candidate answers derived from contexts, considering an answer correct if its normalized form matches any of the golden answers.
Subsequently, the context-conflicting datasets are composed of samples for which only one of the two types of contexts, either generated or retrieved, yields the correct answer, thereby ensuring the exclusivity. Notably, each dataset comprises two distinct subsets: AIG, consisting of samples with correct answers only in the generated context; and AIR comprising samples with correct answers only in the retrieved context.
@@ 3.2 Statistics of Datasets
Table 1: Dataset size across LLMs. “Avg. Prop.” shows average proportions of subsets to original datasets.
@@
We initially adopt a typical and simple setting in which an LLM serves as both the generator and reader. Table 1 provides statistics for the constructed subsets corresponding to various LLMs, including GPT-4 (gpt-4-0613), GPT-3.5 (gpt-3.5-turbo-0613), Llama2-7b/13b (Llama2-7b/13b-chat (Touvron et al., 2023)). The statistics show that the context-conflicting subsets form a substantial part of the datasets, underscoring the need to investigate how LLMs integrate these distinct contexts. Notably, GPT-4 has fewer conflicting instances than other LLMs, because of its higher efficacy in answering questions using either solely retrieved or generated contexts.
Section 4.2 also explores a more complex scenario in which the generator and reader are distinct LLMs, with the statistics shown in Appendix A.2.
@@ 3.3 Evaluation Metric
The metric $\mathrm{DiffGR}$, ranging from $[-1,1]$, quantifies the extent of LLMs’ tendency to rely on generated contexts over retrieved contexts. Using AIR as an example, where all correct answers come from retrieved contexts, an ideal $\mathrm{DiffGR}$ is $-1$, i.e., LLMs should always rely on retrieved contexts.
@@ 4 How LLMs Merge Contexts?
This section conducts experiments on the constructed datasets to investigate the merging mechanism of the LLMs in hybrid approaches. We first consider a typical setting where the generator and reader share a single LLM, to explore how LLMs merge retrieved and self-generated contexts (§4.1). Then, we extend our experiments to include more flexible combinations of generator and reader (§4.2), and various retrieval models (§4.3), to investigate their effects.
@@ 4.1 LLMs Prefer Self-Generated Contexts
Our preliminary experiments, in which a single LLM serves both as generator and reader, are designed to explore how LLMs integrate information from retrieved and self-generated contexts. The LLMs under evaluation are tasked with answering questions using both types of contexts on their corresponding CC datasets. In all experiments, we employ a randomized input sequence of contexts to mitigate the influence of order, which is further discussed in Appendix B.3.
We begin our analysis by examining LLMs’ QA performance on CC datasets to reveal how well can LLMs utilize both types of contexts. Table 2 presents the Exact Match scores (Yu et al., 2022) across various LLMs. Surprisingly, LLMs demonstrate significantly low performance ($\leq 22.16\%$) on AIR subsets, despite the fact that the retrieved context alone consistently yields the correct answer on these subsets. In contrast, LLMs exhibit strong performance on AIG subsets (most near $90\%$). Overall, all LLMs exhibit a significant performance gap between AIR and AIG datasets, with a pronounced decline in performance when the correct answers come from retrieved contexts.
@@
To further reveal LLMs’ behavior underlying the QA performance, we trace the source contexts of LLMs’ answers using the proposed $\mathrm{DiffGR}$ metric. An ideal LLM should always rely on retrieved contexts on AIR subsets ($\mathrm{DiffGR=-1}$), and always rely on generated contexts on AIG subsets ($\mathrm{DiffGR=1}$). Contrary to expectations, Figure 5 illustrates that LLMs fail to identify the correct information and consistently tend to rely on generated contexts on both AIG and AIR subsets. This result indicates a pronounced bias in LLMs to favor generated contexts, even when they provide incorrect information. This bias leads to the insufficient utilization of retrieved contexts mentioned above and highlights a critical challenge for existing LLMs in effectively merging generated and retrieved contexts. As the bias on AIR subsets has a more direct impact on the performance, the following experiments and analysis will focus on the biases on these subsets to conserve space. Results on the AIG subsets can be found in Appendix B.1.
@@ 4.2 LLMs Broadly Prefer Generated Contexts
The above experiments reveal the bias in LLMs to favor the self-generated context. A question emerges: Do LLMs also prefer contexts generated by other LLMs? To investigate this question, this section extends the experiments to more flexible combinations of generators and readers. This setting is also of practical significance, as recent works have explored the decoupling of generators and readers to achieve modularization of knowledge (Luo et al., 2023; Feng et al., 2023).
We construct context-conflicting datasets for each (generator, reader) pair respectively. The statistics of these datasets are shown in Section A.2. Based on these datasets, we then compute $\mathrm{DiffGR}$ metric to examine biases across various (generator, reader) pairs, as shown in Figure 6, and observe two notable insights: (i) LLMs are also biased towards contexts generated by other LLMs.This suggests that such bias in LLMs is widespread and not limited to self-generated contexts. (ii) LLMs usually exhibit a stronger bias to contexts generated by themselves.The sole exception is Llama2-7b, which shows the strongest bias when paired with Llama2-13b as its generator. This phenomenon likely results from their highly similar model structures and training processes (Touvron et al., 2023).